Solve the linear system: 3x+y=5
2x+3y=8
Show all work.

Answer by stanbon(75887)

Solve the linear system:
3x+y=5
2x+3y=8
----
Multiply thru the top equation by 3:
9x + 3y = 15
2x + 3y = 8
---------
Subtract and solve for "x":
7x = 7
x = 1
------
Solve for "y":
3x + y = 5
3*1 + y = 5
y = 2
=======================
cheers,
Stan H.
=======================